the perimeter of an equilateral triangle is 39 cm. find the length of an altitude of the triangle
Answer by stanbon(57222)

the perimeter of an equilateral triangle is 39 cm. find the length of an altitude of the triangle
---------
Draw a picture of an equilateral triangle.
Draw an altitude (which is a perpendicular bisector of the base).
---------
Use Pythagoras to find the altitude:
h^2 + (39/2)^2 = 39^2
---
h^2 = (3/4)39^2
---
h = (39/2)sqrt(3) cm
========================
Cheers,
Stan H.
========================